Before conference swim championships, Suzanne’s best time in the 200 yard butterfly race was 2 minutes, 23 seconds. In the conference championships, she won 10th place by swimming the 200 yard butterfly in 2 minutes, 15 seconds. What per cent decrease is that?
A.3.6% B.5.6% C.5.9% D.6.2% E.36.9%

Answer by Theo(3458)

2 minutes 23 seconds is equivalent to 143 seconds.

2 minutes 15 seconds is equivalent to 135 seconds.

she lowered her time by 8 seconds (135 + 8 = 143).

the percent decrease equals the difference divided by the first time, with the result of that operation being multiplied by 100%. This becomes:

8/143 = .05594405594 * 100% = 5.594405594 percent.

that's pretty close to selection B of 5.6%